OpenStudy (hjkhjkhjkhjk):

Solve: log x− log (2x − 1) = 0

OpenStudy (hjkhjkhjkhjk):

Select one: a. 10/19 b. No solution c. 0 d. 1

jhonyy9 (jhonyy9):

use formula log a - log b = log a/b

OpenStudy (hjkhjkhjkhjk):

This would be base 10, correct?

jhonyy9 (jhonyy9):

yes but just add to both sides log(2x-1) and so what will get

jhonyy9 (jhonyy9):

this is very easy to solve it

OpenStudy (mathmale):

Kindly show your work if you want meaningful feedback on it.

OpenStudy (mathmale):

No change of base is needed here. Where did that suggestion come from?

OpenStudy (hybrik):

log x -log(2x-1)=0 x-2x-1=0 (i did 10^ (log x)-10^ log(2x-1)=10^ 0)

jhonyy9 (jhonyy9):

how you think thi 10^0 = 1 not zero ?

OpenStudy (hjkhjkhjkhjk):

I found the answer, thank you.

OpenStudy (hybrik):

oh yeah, i forgot i didnt change it, np

OpenStudy (mathmale):

Note 1: "log" implies that the base is 10. No need to actually write the "10." Note 2: "log x− log (2x − 1)" is the difference of two logs; according to the "division rule" for logs, log x− log (2x − 1) = \[\log\frac{ x }{ (2x-1) }\]

OpenStudy (mathmale):

Please solve the original problem for x. Hint: the log of what number is zero?

jhonyy9 (jhonyy9):

@mathmale sorry but my idea may be that log x - log (2x-1)= 0 log x = log (2x-1) so x = 2x-1 1 = x your opinion please ?

jhonyy9 (jhonyy9):

ok. hjkhjkhjk ? choice d. is right sure

jhonyy9 (jhonyy9):

this is what you got too :?

jhonyy9 (jhonyy9):

@IrishBoy123

jhonyy9 (jhonyy9):

@ganeshie8 @TheSmartOne

OpenStudy (irishboy123):

sure! think that is the way mm was going